# MaskINT: Video Editing via Interpolative Non-autoregressive Masked Transformers

### v1

@@ Abstract
Recent advances in generative AI have significantly enhanced image and video editing, particularly in the context of text prompt control. State-of-the-art approaches predominantly rely on diffusion models to accomplish these tasks. However, the computational demands of diffusion-based methods are substantial, often necessitating large-scale paired datasets for training, and therefore challenging the deployment in practical applications. This study addresses this challenge by breaking down the text-based video editing process into two separate stages. In the first stage, we leverage an existing text-to-image diffusion model to simultaneously edit a few keyframes without additional fine-tuning. In the second stage, we introduce an efficient model called MaskINT, which is built on non-autoregressive masked generative transformers and specializes in frame interpolation between the keyframes, benefiting from structural guidance provided by intermediate frames. Our comprehensive set of experiments illustrates the efficacy and efficiency of MaskINT when compared to other diffusion-based methodologies. This research offers a practical solution for text-based video editing and showcases the potential of non-autoregressive masked generative transformers in this domain.
@@ 1 Introduction
Recently, numerous studies have showcased the impressive capabilities of diffusion models [18] in the domain of text-to-image and text-to-video generation [37, 38, 3, 40]. Later on, built upon Stable Diffusion (SD) [37], several works have achieved remarkable success in the realm of text-based image editing [52, 4, 43]. When extending to text-based video editing, existing works can be mainly divided into two ways: One is to train diffusion models with temporal modules on paired text-video datasets [19, 13]. However, due to the lack of extensive text-to-video datasets, these works typically struggle to achieve the same level of editing expertise seen in the realm of image editing. The other involves leveraging a pre-trained text-to-image diffusion models in a training-free techniques [15, 34, 25, 53]. These works usually extend the self-attention across all frames to achieve an overall temporal consistency. However, this attention-based temporal constraint remains implicit and suboptimal. Moreover, while diffusion-based techniques are capable of producing high-fidelity videos, the use of diffusion models to produce all video frames proves to be a highly time-consuming process. The integration of the global attention across all frames in these video editing methods further extends the processing time, rendering them less practical for real-world applications.
Meanwhile, studies indicate that non-autoregressive masked generative transformers [7, 8, 51, 16] can attain similar levels of performance in generating images or videos compared to diffusion-based methods, while bring significant efficiency [8]. These works first tokenize image or videos into a sequence of discrete tokens [12], and then train transformers [44, 10] with masked token modeling to predict these tokens. During the inference time, they employ non-autoregressive decoding, which generates all tokens in parallel and iteratively refine predictions in a few steps. Nonetheless, extending these techniques to perform global editing tasks, such as stylization, presents a formidable challenge. This type of task requires the replacement of nearly all tokens with new ones, as opposed to merely modifying tokens within a localized region.
@@
In this paper, we disentangle text-based video editing into two separate stages. In the first stage, we utilize existing text-based image editing models to jointly edit only two keyframes (i.e., the initial and last frames) from the video, guided by the provided text prompt. In the second stage, we propose a novel Interpolative Non-autoregressive generative Transformers (MaskINT), which performs structure-aware frame interpolation by leveraging both the color information of the initial and final frames, as well as structural cues like edge or depth maps from intermediate frames. Through disentanglement of keyframe editing and frame interpolation into separate stages, our pipeline eliminates the requirement for paired video datasets during training, thereby enabling us to train the MaskINT using video-only datasets. Furthermore, thanks to its non-autoregressive decoding, MaskINT significantly accelerates the generation of intermediate frames compared to using diffusion models for this purpose. We show that our method balances the trade off between quality and efficiency, offering comparable performance with existing diffusion methods yet taking much less time in generation.
@@
  • We propose to disentangle the text-based video editing into a two stage pipeline, that involves keyframes joint editing using existing image diffusion model and structure-aware frame interpolation with masked generative transformers trained on video only datasets.
  • We propose MaskINT to perform structure-aware frame interpolation, which is the pioneer work that explicitly introduces structure control into non-autoregressive generative transformers.
  • Experimental results demonstrate that our method achieves comparable performance with diffusion methods in terms of temporal consistency and alignment with text prompts, while providing 5-7 times faster inference times.
@@ 2 Related Work
Following GPT [5], many pioneer works [12, 50, 26, 14, 20] tokenize images/videos into discrete tokens, and train Autoregressive Generative Transformers to perform image/video generation, where tokens are generated sequentially based on previous output. However, these autoregressive methods become exceedingly time-consuming when the length of the token sequence increases. Recently, Non-autoregressive Generative Transformers, capable of simultaneously generating all tokens in parallel, have emerged as efficient solutions [7, 51]. Specifically, MaskGiT [7] first shows the capability and efficiency of this technique in image generation. It can be seamlessly extended to tasks like inpainting and extrapolation by applying various initial mask constraints. Muse [8] achieves state-of-the-art performance in text-to-image generation by training on large-scale text-image datasets and brings significantly efficiency improvement. StyleDrop [41] further finetunes Muse with human feedback to perform text-to-image generation guided with a reference style image. Furthermore, MaskSketch [1] introduces implicit structural guidance into MaskGiT by calculating the similarity of attention maps in the sampling step. Nevertheless, this implicit structure condition is suboptimal.
In video generation, MaskViT [16] employ 2D tokenizer and trains a bidirectional window transformer to perform frame prediction. Phenaki [45] trains a masked transformer to generate short video clips condition on text prompt and extends it to arbitrary long video with different prompts in an autoregressive way. MAGVIT [51] utilizes 3D tokenizer to quantize videos and trains a single model to perform multiple video generation tasks such as inpainting, outpainting, frame interpolation, etc. However, to the best of our knowledge, there is currently no existing literature in the field of text-based video editing utilizing masked generative transformers. Besides, there is a notable absence of research that delves into explicit structural control within this area.
@@
Leveraging the advancements of SD [37], numerous studies have achieved significant success in the field of text-based image editing [52, 30, 43, 24, 17, 54, 32, 29, 9]. For example, ControlNet [52], T2I-Adapter [30], and Composer [21] finetune SD with spatial condition such as depth maps and edge maps, enabling text-to-image synthesis with the same structure as the input image. The PNP [43] incorporates DDIM inversion features [42] from the input image into the text-to-image generation process alongside SD, enabling image editing without the necessity of additional training or fine-tuning. Instructpix2pix [4] trains a conditional diffusion model for text-guided image editing using synthetic paired examples, which avoid the tedious inversion. Nevertheless, employing these methods on each video frame independently often leads to inconsistencies and flickering.
@@
Recently, diffusion models also dominate the field of video generation and video editing [40, 19]. For example, Gen-1 [13] trains a video diffusion models on paired text-video datasets to generate videos with both depth map and text prompts. Meanwhile, several works utilize pre-trained image diffusion models to achieve video editing in a training-free way [47, 49, 34, 25, 53, 46, 6]. To enable a cohesive global appearance among edited frames, a common approach in these studies involves extending the attention module of SD to encompass multiple frames and conducting cross-frame attention. In detail, Text2Video-Zero [25] performs cross-frame attention of each frame on the first frame to preserve appearance consistency. ControlVideo [53] extends ControlNet with fully cross-frame attention to joint edit all frames and further improves the performance with interleaved-frame smoother. TokenFlow [15] enhance PNP [43] with extended-attention to jointly edit a few keyframes at each denoising step and propagate them throughout the video based on nearest-neighbor field. On the contrary, we only utilize existing image diffusion models to edit two keyframes, rather than all video frames.
@@
VFI aims to generate intermediate images between a pair of frames, which can be applied to creating slow-motion videos and enhancing refresh rate. Advanced methods typically entail estimating dense motions between frames, like optical flow, and subsequently warping the provided frames to generate intermediate ones [31, 23, 39, 28, 36, 22]. However, these methods are most effective with simple or monotonous motion. Thus, they cannot be directly applied to the second stage. In our work, we perform frame interpolation by incorporating additional structural signals.
@@ 3.1 Masked Generative Transformers
Masked generative transformers [7] follow a two-stage pipeline. In the first stage, an image is quantized into a sequence of discrete tokens via a Vector-Quantized (VQ) auto-encoder [12]. In detail, given an image $\textbf{I}\in\mathbb{R}^{H\times W\times 3}$, an encoder $\mathcal{E}$ encodes it into a series of latent vectors and discretize them through a nearest neighbour look up in a codebook of quantized embeddings with size $M$. To this end, an image can be represented with a sequence of codebook’s indices $\mathbf{Z}=[z_{i}]_{i=1}^{h\times w},z_{i}\in\{1,2,...,M\}$, where $h$ and $w$ is the resolution of latent features. A decoder $\mathcal{D}$ can reconstruct the indices back to image $\mathcal{D}(\mathbf{Z})\approx\mathbf{I}$. In the second stage, a bidirectional transformer model [44] is learned with Masked Token Modeling (MTM). Specifically, during training, a random mask ratio $r\in(0,1)$ is selected and $\lfloor\gamma(r)\cdot h\times w\rfloor$ tokens in $Z$ are replaced with a special [MASK] token, where $\gamma(r)$ is a mask scheduling function [7]. We denote the corrupted sequence with masked tokens as $\mathbf{\bar{Z}}$ and conditions such as class labels or text prompt as $\mathbf{c}$. Given the training dataset $\mathbb{D}$, a BERT [10] parameterized by $\Phi$ is learned to minimize the cross-entropy loss between the predicted and the ground truth token at each masked position:
@@
During inference time, non-autoregressive decoding is applied to generate images. Specifically, given the conditions $\mathbf{c}$, all tokens are initialized as [MASK] tokens. At step $k$, all tokens are predicted in parallel while only tokens with the highest prediction scores are kept. The remaining tokens with least prediction scores are masked out and regenerated in the next iteration. The mask ratio is determined by $\gamma(\frac{k}{K})$ at step $k$, where $K$ is the total number of iteration steps.
@@ 3.2 ControlNet
Denoising Diffusion Probabilistic Models (DDPM) [18] generate images through a progressive noise removal process applied to an initial Gaussian noisy image, carried out over a span of $T(>>K)$ time steps. To enable efficient high-resolution image generation, Latent Diffusion models [37] operates the diffusion process in the latent space of an autoencoder. First, an encoder $\mathcal{E^{\prime}}$ compresses an image $\mathbf{I}$ to a low-resolution latent code $x=\mathcal{E^{\prime}}(\mathbf{I})\in\mathbb{R}^{h\times w\times c}$. Second, a U-Net $\epsilon_{\theta}$ with attention modules [44] is trained to remove the noise with loss function:
$\mathcal{L}_{LDM}=\mathbb{E}_{x_{0},\epsilon\sim N(0,I),t\sim}\|\epsilon-\epsilon_{\theta}(x_{t},t,\tau)\|_{2}^{2},$ (2)
where $\tau$ is the text prompt and $x_{t}$ is the noisy latent sample at timestep $t$. Stable Diffusion is trained on datasets with billion scale text-image pairs, which serves as the foundation model of many generation tasks.
@@
In practice, it’s challenging to use text prompt to describe the layout of generated image. Furthermore, ControlNet [52] is proposed to provide spatial layout conditions such as edge map, depth map, and human poses. In detail, ControlNet train the same U-Net architecture as SD and finetune it with specific conditions. We denote ControlNet as $\epsilon_{\theta}^{\prime}(x_{t},t,\tau,s)$, where $s$ is the spatial layout condition.
@@ 4.1 Overview
Fig. 2 shows an overview of our framework. We disentangle the video editing task into keyframe joint editing stage and structure-aware frame interpolation stage. Specifically, given a video clip with $N$ frames $\{\mathbf{I}^{i}\}_{i=1}^{N}$, in the first stage, with the input text prompt $\mathbf{\tau}$, we simultaneously edit two keyframes, i.e., the initial frame $\mathbf{I}^{1}$ and last frame $\mathbf{I}^{N}$, using existing image-editing model $g(\cdot)$ that requires no additional tuning. This frame based joint editing module provides high-quality coherent edited frames $\mathbf{{}^{*}I}^{1},\mathbf{{}^{*}I}^{N}=g(\mathbf{I}^{1},\mathbf{I}^{N},\mathbf{\tau})$ and is highly efficient on a pair of frame. In the second stage, we propose MaskINT to perform structure-aware frame interpolation via non-autoregressive transformers. The entire edited video frames are generated by $\{\mathbf{{}^{*}I}^{i}\}_{i=1}^{N}=f_{\Phi}(\mathbf{{}^{*}I}^{1},\mathbf{{}^{*}I}^{N},\{\mathbf{S}^{i}\}_{i=1}^{N})$, where $\mathbf{S}^{i}\in[0,1]^{H\times W\times 1}$ is the structural condition (i.e., the HED edge map [48]). Our MaskINT is trained with masked token modeling (MTM) on video only datasets, conditioning on the structural signal $\{\mathbf{S}^{i}\}_{i=1}^{N}$ as well as the initial frame $\mathbf{I}^{1}$ and last frame $\mathbf{I}^{N}$.
@@ 4.2 Keyframes Joint Editing
To maintain the structure layout of selected keyframes, we take ControlNet [52] to edit $\mathbf{I}^{1}$ and $\mathbf{I}^{N}$ based on text prompt $\tau$ as well as their edge maps $\mathbf{S}^{1}$ and $\mathbf{S}^{N}$. However, even with the identical noise, applying ControlNet to each keyframes individually, i.e., $\epsilon_{\theta}^{\prime}(x_{t}^{1},t,\tau,\mathbf{S}^{1})$ and $\epsilon_{\theta}^{\prime}(x_{t}^{N},t,\tau,\mathbf{S}^{N})$, cannot guarantee the appearance consistency. To address this issue, following previous work [47, 15, 53], we extend the self-attention blocks to simultaneously process two keyframes. In detail, the self-attention block projects the noisy feature map $x_{t}^{j}$ of $j^{th}$ frame at time step $t$ into query $\mathbf{Q}^{j}$, key $\mathbf{K}^{j}$, value $\mathbf{V}^{j}$ in the original U-Net of SD. We extend the self-attention block to perform attention across all selected keyframes by concatenating their keys and values and calculate the attention by
$\text{Softmax}(\dfrac{\mathbf{Q}^{j}[\mathbf{K}^{1},\mathbf{K}^{N}]^{\mathbf{T}}}{\sqrt{c}})[\mathbf{V}^{1},\mathbf{V}^{N}]$ (3)
Note that although only two frames were used in Eq.3, this joint editing can seamlessly generalize to any number of frames, but at the cost of longer processing time and huge demand in resources.
4.3 MaskINT

### v2

@@ Abstract
Recent advances in generative AI have significantly enhanced image and video editing, particularly in the context of text prompt control. State-of-the-art approaches predominantly rely on diffusion models to accomplish these tasks. However, the computational demands of diffusion-based methods are substantial, often necessitating large-scale paired datasets for training, and therefore challenging the deployment in real applications. To address these issues, this paper breaks down the text-based video editing task into two stages. First, we leverage an pre-trained text-to-image diffusion model to simultaneously edit few keyframes in an zero-shot way. Second, we introduce an efficient model called MaskINT, which is built on non-autoregressive masked generative transformers and specializes in frame interpolation between the edited keyframes, using the structural guidance from intermediate frames. Experimental results suggest that our MaskINT achieves comparable performance with diffusion-based methodologies, while significantly improve the inference time. This research offers a practical solution for text-based video editing and showcases the potential of non-autoregressive masked generative transformers in this domain.
@@ 1 Introduction
Recently, numerous studies have showcased the impressive capabilities of diffusion models [22] in image generation and editing with text prompts [44, 13, 39, 45, 61, 6, 51]. When extending these foundation text-to-image (T2I) models to video generation and editing, a challenge is to maintain the temporal consistency. Existing solutions can be mainly divided into two ways: One is to finetune T2I models with additional temporal modules on paired text-video datasets [23, 17, 9, 32, 5]. However, the lack of extensive paired text-video datasets hinders these works from attaining the same level of performance as seen in image editing. The other involves leveraging a T2I models in a zero-shot manner [19, 41, 29, 8, 62]. These works find that extending the spatial self-attention across frames can produce consistent content across frames [56]. However, extended attention can lead to inconsistent editing of details due to its implicit temporal constraints. Moreover, while these techniques can generate high-fidelity videos, using cumbersome diffusion models to generate all frames is time-consuming. The integration of the extended spatial-temporal attention further exacerbates the running time, making them less practical for real-world applications. This raises the question: Is it necessary to utilize tedious diffusion models to generate all video frames?
@@
In this paper, we propose MaskINT, an efficient prompt-based video editing framework designed to generate high-quality videos swiftly (Fig. 1). MaskINT disentangles the task into two separate stages. In the first stage, we utilize pre-trained T2I models with extended attention to jointly edit only two keyframes (i.e., the initial and last frames) from the video clip, guided by the provided text prompt. In the second stage, we introduce a novel structure-aware frame interpolation module based on non-autoregressive generative Transformers [10, 11, 60], which generates all intermediate frames in parallel with structural cues and iteratively refine predictions in a few steps. Our framework offers several advantages: First, through the disentanglement of frame editing and interpolation, MaskINT eliminates the requirement for paired text-video datasets during training, thereby enabling us to train the MaskINT using large-scale video-only datasets. Second, the usage of non-autoregressive generation significantly accelerates the processing time. Experimental results indicate that MaskINT achieves comparable performance with pure diffusion methods in terms of temporal consistency and alignment with text prompts, while provides 5-7 times faster inference time. Our work not only provides a practical solution for text-based video editing in terms of the trade-off between quality and efficiency, but also highlights the potential of masked generative transformers in this domain.
@@
  • We introduce MaskINT, a two-stage pipeline for text-based video editing, which involves keyframes joint editing and structure-aware frame interpolation, eliminating the need for paired text-video datasets.
  • The proposed structure-aware frame interpolation is the pioneer work that explicitly introduces structure control into non-autoregressive generative transformers.
  • Experimental results demonstrate that MaskINT achieves comparable performance with diffusion methods, while significantly reducing the inference time.
@@ 2 Related Work
Following GPT [7], previous works [16, 59, 31, 18, 24] first tokenize images/videos into discrete tokens, and train Autoregressive Generative Transformers to perform image/video generation, where tokens are generated sequentially based on previous outputs. However, these methods become exceedingly time-consuming when the length of the token sequence increases.
Recently, Non-autoregressive Generative Transformers have emerged as efficient solutions [10, 11, 60, 20, 54]. These works train bidirectional transformers [14] with masked token modeling (MTM). During inference, all discrete tokens are generated in parallel and refined iteratively in a few steps, known as non-autoregressive decoding. Specifically, in image generation, MaskGiT [10] first shows the capability and efficiency of non-autoregressive transformers. It can be seamlessly extended to tasks like inpainting and outpainting by applying various initial mask constraints. Muse [11] achieves state-of-the-art performance in text-to-image generation by training on large-scale text-image datasets and brings significantly efficiency improvement. StyleDrop [48] further finetunes Muse with human feedback to perform text-to-image generation guided with a reference style image. MaskSketch [3] introduces implicit structural guidance into MaskGiT by calculating the similarity of attention maps in the sampling step. In video generation, MaskViT [20] employ 2D tokenizer and trains a bidirectional window transformer to perform frame prediction. Phenaki [54] trains a masked transformer to generate short video clips condition on text prompt and extends it to arbitrary long video with different prompts in an autoregressive way. MAGVIT [60] utilizes 3D tokenizer to quantize videos and trains a single model to perform multiple video generation tasks such as inpainting, outpainting, frame interpolation, etc. However, to the best of our knowledge, there is currently no existing literature in the field of text-based video editing utilizing masked generative transformers. Besides, there is a notable absence of research that delves into explicit structural control within this area.
@@
Built upon the latent diffusion models (LDM) [44], numerous studies have achieved significant success in text-based image editing [61, 36, 51, 28, 21, 63, 38, 35, 12]. For example, ControlNet [61], T2I-Adapter [36], and Composer [25] finetune LDM with spatial condition such as depth maps and edge maps, enabling text-to-image synthesis with the same structure as the input image. The PNP [51] incorporates DDIM inversion features [49] from the input image into the T2I generation process, enabling zero-shot image editing with pre-trained T2I models. InstructPix2Pix [6] trains a conditional diffusion model for text-guided image editing using synthetic paired examples, which avoid the tedious inversion. Nevertheless, applying these T2I models on each frame independently usually leads to inconsistencies and flickering.
@@
Diffusion models also dominate the field of video generation and video editing [47, 23, 9, 56]. Due to the lack of billion-scale paired video datasets, these works usually finetune existing T2I foundation models [44] with additional temporal modules on paired text-video datasets [17] or overfit a single video [56]. Meanwhile, several works explore using T2I models in a zero-shot manner for video editing [56, 58, 41, 29, 62, 55, 8]. To enable a cohesive appearance, a common approach in these studies involves extending the spatial self-attention module to spatial-temporal attention module and conducting cross-frame attention. In detail, Text2Video-Zero [29] performs cross-frame attention of each frame on the first frame to preserve appearance consistency. ControlVideo [62] extends ControlNet [61] with fully cross-frame attention to joint edit all frames and further improves the performance with interleaved-frame smoother. TokenFlow [19] enhance PNP [51] with extended-attention to jointly edit a few keyframes at each denoising step and propagate the edit frames to the entire video. However, conducting hundreds of denoising steps with extended attention across all frames is time-consuming. In contrast, we only utilize diffusion models to edit a few keyframes, rather than all frames.
@@
VFI aims to generate intermediate images between a pair of frames, which can be applied to creating slow-motion videos and enhancing refresh rate. Advanced methods typically estimate dense motions between frames, like optical flow, and subsequently warping the provided frames to generate intermediate ones [37, 27, 46, 34, 43, 26]. However, these regression methods are most effective with simple or monotonous motion. In our work, we design a generative model to perform frame interpolation with additional structural signals, enabling interpolation with large motion.
@@ 3.1 Masked Generative Transformers
Masked generative transformers [10] follow a two-stage pipeline. In the first stage, an image is quantized into a sequence of discrete tokens via a Vector-Quantized (VQ) auto-encoder [16]. In detail, given an image $\textbf{I}\in\mathbb{R}^{H\times W\times 3}$, an encoder $\mathcal{E}$ encodes it into a series of latent vectors, which are then discretized via nearest neighbor lookup in a codebook of quantized embeddings with size $M$. Thus, an image can be represented with a sequence of codebook’s indices $\mathbf{Z}=[z_{i}]_{i=1}^{h\times w},z_{i}\in\{1,2,...,M\}$, where $h$ and $w$ is the resolution of latent features. A decoder $\mathcal{D}$ can reconstruct the indices back to image $\mathcal{D}(\mathbf{Z})\approx\mathbf{I}$. In the second stage, a bidirectional transformer model [53] is learned with Masked Token Modeling (MTM). Specifically, during training, a random mask ratio $r\in(0,1)$ is selected and $\lfloor\gamma(r)\cdot h\times w\rfloor$ tokens in $Z$ are replaced with a special [MASK] token, where $\gamma(r)$ is a mask scheduling function [10]. We denote the corrupted sequence with masked tokens as $\mathbf{\bar{Z}}$ and conditions such as class labels or text prompt as $\mathbf{c}$. Given the training dataset $\mathbb{D}$, a BERT [14] parameterized by $\Phi$ is learned to minimize the cross-entropy loss between the predicted and the ground truth token at each masked position:
@@
During inference, non-autoregressive decoding is applied to generate images. Specifically, all tokens of an image or video are initialized as [MASK] tokens. At step $k$, all tokens are predicted in parallel while only tokens with the highest prediction scores are kept. The remaining tokens with least prediction scores are masked out and regenerated in the next iteration. The mask ratio is determined by $\gamma(\frac{k}{K})$ at step $k$, where $K$ is the total number of iteration steps.
@@ 3.2 ControlNet
Denoising Diffusion Probabilistic Models [22] generate images through a progressive noise removal process applied to an initial Gaussian noisy image, carried out over a span of $T$ time steps. To enable efficient high-resolution image generation, LDM [44] operates the diffusion process in the latent space of an autoencoder. In detail, an encoder $\mathcal{E^{\prime}}$ first compresses an image $\mathbf{I}$ to a low-resolution latent code $x=\mathcal{E^{\prime}}(\mathbf{I})\in\mathbb{R}^{h\times w\times c}$. Second, a U-Net $\epsilon_{\theta}$ with attention modules [53] is trained to remove the noise with loss
$\mathcal{L}_{LDM}=\mathbb{E}_{x_{0},\epsilon\sim N(0,I),t\sim}\|\epsilon-\epsilon_{\theta}(x_{t},t,\tau)\|_{2}^{2}$ (2)
, where $\tau$ is the text prompt and $x_{t}$ is the noisy latent sample at timestep $t$.
@@
Since accurately describing the layout of generated images using only text prompts is challenging, ControlNet [61] extends LDM by incorporating spatial layout conditions such as edge maps, depth maps, and human poses. In detail, ControlNet train the same U-Net architecture as LDM and finetune it with additional conditions. We denote it as $\epsilon_{\theta}^{\prime}(x_{t},t,\tau,s)$, where $s$ is the spatial layout.
@@ 4.1 Overview
Fig. 2 shows an overview of our MaskINT. MaskINT disentangle the video editing task into keyframe editing stage and structure-aware frame interpolation stage. Specifically, given a video clip with $N$ frames $\{\mathbf{I}^{i}\}_{i=1}^{N}$, in the first stage, with the input text prompt $\mathbf{\tau}$, we simultaneously edit two keyframes, i.e., the initial frame $\mathbf{I}^{1}$ and last frame $\mathbf{I}^{N}$, using existing image-editing model $g(\cdot)$ in a zero-shot way. This process results in the creation of high-quality coherent edited frames $\mathbf{{}^{*}I}^{1},\mathbf{{}^{*}I}^{N}=g(\mathbf{I}^{1},\mathbf{I}^{N},\mathbf{\tau})$. In the second stage, MaskINT performs structure-aware frame interpolation via non-autoregressive transformers. The entire edited video frames are generated by $\{\mathbf{{}^{*}I}^{i}\}_{i=1}^{N}=f_{\Phi}(\mathbf{{}^{*}I}^{1},\mathbf{{}^{*}I}^{N},\{\mathbf{S}^{i}\}_{i=1}^{N})$, where $\mathbf{S}^{i}\in[0,1]^{H\times W\times 1}$ is the structural condition (i.e., the HED edge map [57]). The structure-aware interpolation module is trained with masked token modeling (MTM) on video only datasets, conditioning on the structural signal $\{\mathbf{S}^{i}\}_{i=1}^{N}$ as well as the initial frame $\mathbf{I}^{1}$ and last frame $\mathbf{I}^{N}$.
@@ 4.2 Keyframes Joint Editing
To maintain the structure layout, we take ControlNet to edit $\mathbf{I}^{1}$ and $\mathbf{I}^{N}$ based on text prompt $\tau$ as well as their edge maps $\mathbf{S}^{1}$ and $\mathbf{S}^{N}$. However, even with the identical noise, applying ControlNet to each keyframes individually cannot guarantee the appearance consistency. To address this issue, following previous work [56, 19, 62], we extend the self-attention to simultaneously process two keyframes. In detail, the self-attention module projects the noisy feature map $x_{t}^{j}$ of $j^{th}$ frame at time step $t$ into query $\mathbf{Q}^{j}$, key $\mathbf{K}^{j}$, value $\mathbf{V}^{j}$ in the original U-Net. We extend the self-attention block to perform attention across all selected keyframes by concatenating their keys and values:
$\text{Softmax}(\dfrac{\mathbf{Q}^{j}[\mathbf{K}^{1},\mathbf{K}^{N}]^{\mathbf{T}}}{\sqrt{c}})[\mathbf{V}^{1},\mathbf{V}^{N}].$ (3)
Since each frame queries all frames and aggregates information from all of them, the appearance consistency can be achieved. Although only two frames were used in Eq. 3, this joint editing can be seamlessly extended to any number of frames, but at the cost of increased processing time and significant GPU memory demand.
4.3 Structure-Aware Frame Interpolation